# “Stones from Other Hills can Polish Jade”: Zero-shot Anomaly Image Synthesis via Cross-domain Anomaly Injection

### v1

@@ 1 Introduction
To ease above problems, recent years have witnessed a surging interest in few-shot/zero-shot anomaly synthesis (FSAS/ZSAS), which aims to synthesize pseudo anomaly images for a specific industrial domain with few or zero domain-specific anomaly. A fair number of FSAS/ZSAS solutions have been proposed (reviewed in Sec. 2), while pseudo anomalies gradually occupy an indispensable role in many recent IAD solutions [zavrtanik2021draem, li2021cutpaste, schluter2022natural, zhang2024realnet, hu2024anomalydiffusion]. Since FSAS still requires accessing a few domain-specific anomalies in advance, FSAS suffers from the uncertainty or even unavailability of domain-specific anomalies. By contrast, ZSAS makes no assumption on the availability of domain-specific anomalies, thus enjoying better applicability to more diverse domains. Hence, we will focus on ZSAS in this paper. To maximize synthetic anomalies’ utility and facilitate them to be incorporated into IAD, a natural goal is to realize both authentic and simple ZSAS. Unfortunately, we observe that existing ZSAS solutions find it hard to balance authenticity and simplicity. Specifically, existing ZSAS solutions can be categorized into model-free solutions and model-based solutions: Model-free solutions (e.g. CutPaste [li2021cutpaste], DRAEM [zavrtanik2021draem]) play a dominant role in ZSAS. They sample patterns from given normal images or external datasets that are irrelevant to IAD, and directly inject those patterns into the target normal image to synthesize anomalies. In this way, such solutions enjoy a simple use without training an extra model. However, pseudo anomalies synthesized by model-free solutions typically lack authenticity, because their pseudo anomaly patterns are irrelevant to real anomalies. Besides, their authenticity is further diminished by injection artifacts (e.g. pixel discontinuity). By contrast, recent model-based solutions like RealNet [zhang2024realnet] place more emphasis on authenticity. They promote authenticity by training a deep generative model like Generative Adversarial Networks (GANs) [goodfellow2020generative] or Diffusion Models (DMs) [song2019generative], while the core idea is to re-direct the model’s generative process by some external factors (e.g. carefully-crafted perturbations) to generate pseudo anomalies. Although authenticity is improved, such solutions involve a cumbersome training process that requires heavy calculations and extra hardware. Meanwhile, re-directing the generative process is not straightforward, which often requires delicate model design and tuning. Therefore, model-based solutions lack simplicity due to their training-based nature, and their high cost limits their application to practical IAD.
@@
  • We build the very first domain-agnostic anomaly dataset within our best knowledge, so as to supply CAI with rich real anomaly patterns from abundant cross-domain anomalies. It also benefits future ZSAS/IAD research.
  • We propose a CAI-guided Diffusion Mechanism, which can utilize CAI’s authentic pseudo anomaly images to perform unlimited anomaly synthesis. It enables us to further break the quantity limit of real anomalies.
@@ 3.1 Cross-domain Anomaly Injection (CAI)
Given a set of cross-domain anomalies, we randomly sample a cross-domain anomaly image $I_{i}^{(ca)}\in\mathbb{R}^{H_{a}\times W_{a}\times 3}$ as source image, and its binary mask $y_{i}\in\{0,1\}^{H_{a}\times W_{a}}$ that indicates the area of its anomaly pattern $A_{i}=I_{i}^{(ca)}(y_{i})$, where $I_{i}^{(ca)}(y_{i})$ refers to the image pattern indexed by $y_{i}$’s non-zero elements. Then, we fetch a target image $I^{(n)}\in\mathbb{R}^{H_{n}\times W_{n}\times 3}$ from the normal training set of target domain. Finally, CAI intends to inject $A_{i}$ into $I^{(n)}$ to synthesize a pseudo anomaly image $I^{(sa)}$. Concretely, CAI consists of four major stages: Source-target Matching, Multi-scale Anomaly Synthesis, Injection Location Selection and PE-based Anomaly Injection. The procedure of CAI is shown in Fig. 3 and Algo. 1, while we elaborate each stage below:
@@
Source-target Matching. As the source image $I_{i}^{(ca)}$ and target image $I^{(n)}$ are from different domains, real anomaly pattern $A_{i}$ in $I_{i}^{(ca)}$ often requires resizing before injection to fit $I^{(n)}$. In this process, we need to avoid improper resizing, which makes injected anomaly patterns too trivial or too vague. To this end, we propose Source-target Matching (STM). Specifically, we first extract a binary mask to indicate the foreground in $I^{(n)}$: $y^{(n)}=m(I^{(n)})$, where the function $m(\cdot)$ either returns the mask of the object by image matting [gatis_rembg_2022], or return an all-one mask for images without any object (e.g. texture). Afterwards, we calculate the scale of the anomaly pattern in $I_{i}^{(ca)}$ and the foreground in $I^{(n)}$, and define Original Source-Target Ratio (STR), $R$:
@@
Multi-scale Anomaly Synthesis. To promote the diversity in pseudo anomalies’ scale, CAI adopts a Multi-scale Anomaly Synthesis (MAS) strategy. Given the target image, the number of large/medium/small pseudo anomalies, $N_{l}/N_{m}/N_{s}$, are explicitly set for anomaly synthesis. Meanwhile, MAS requires prioritizing synthesis of larger pseudo anomalies under the constraints of STM. For example, MAS prioritizes synthesizing large pseudo anomalies when given a large anomaly pattern, while medium or small pseudo anomalies are produced only when the required number of large pseudo anomaly is reached. Therefore, an anomaly pattern prefers synthesizing a pseudo anomaly of the same scale, but it can also synthesize pseudo anomalies in other scales. In this way, a rational $R^{\prime}$ is selected by STM and MAS (see line 6-23 of Algo. 1). In this paper, we synthesize $N_{l}=4,N_{m}=3,N_{s}=3$ pseudo anomalies with a target normal image, which is a relatively balanced ratio.
@@
where $\partial\Omega(I^{(n)},l_{c},A^{\prime}_{i})$ denotes the border of region $\Omega(I^{(n)},l_{c},A^{\prime}_{i})$. Eq. (3) actually encourages $\Omega(I^{(n)},l_{c},A^{\prime}_{i})$ to share the pattern of the injected anomaly $A^{\prime}_{i}$ by minimizing their difference in gradients, while the constraint of Eq. (3) ensures a smooth transition at the region border. Eq. (3) is equivalent to solving a Poisson partial differential equation that subjects to Dirichlet border condition. Then, we filter out the injected anomaly pattern outside the foreground region by the foreground mask $y^{(n)}$. Considering all inputs, the whole PE-based injection process can be denoted by $I^{(sa)}=PEI(I^{(n)},y^{(n)},l_{c},A^{\prime}_{i})$.
@@
3.2 Domain-agnostic Anomaly Dataset
@@
One of CAI’s foundation is a real anomaly dataset that can provide abundant cross-domain anomalies. However, we notice that existing IAD datasets typically follow a domain-oriented fashion, i.e. they focus on collecting normal/abnormal images under certain domains, and anomalies are mainly categorized by their associated domain. For example, for the popular MvTec-AD dataset [bergmann2019mvtec], its primary categories are like “Bottle/Capsule/Tile…”. Unfortunately, this fashion suffers from two major drawbacks: (1) It hinders the collection of extensive real anomalies, as only anomalies in a certain domain are of interest. (2) It stresses the domain rather than anomalies. Nevertheless, the proposed CAI requires massive cross-domain anomalies, which makes existing IAD datasets short in quantity and diversity of real anomalies. This inspires us to build the first domain-agnostic anomaly dataset, and we show its building procedure in Fig. 4 and detail its key steps below.
@@
Real Anomaly Collection. Real anomalies are actually abundant when the constraint of domain is removed. To collect domain-agnostic real anomalies, we exploit three sources that are all simple and affordable for the general public: (1) Online search: We feed IAD-related textual queries (e.g. “industry scratch”) into the search engine to obtain real anomaly images. (2) Reality photographing: We use frequently-seen portable photographing devices like digital cameras or cellphones to capture real anomalies. (3) Existing datasets: We exploit real anomalies from 9 existing IAD datasets (listed in supplementary material). By virtue of three sources above, we are able to collect about 6000 real anomaly images in a short period ($\sim$ 2 weeks) as the data foundation of our domain-agnostic anomaly dataset. Our anomaly collection is much more extensible when compared with classic IAD datasets, as our dataset welcomes any real anomaly from diverse domains. Note that if an existing dataset is used for evaluation, its data will be strictly excluded from ZSAS to prevent information leakage.
@@
Anomaly Categorization and Annotation. After collecting sufficient real anomalies, we categorize real cross-domain anomalies into several anomaly classes $\{\mathcal{C}_{1},\mathcal{C}_{2},...\}$ in an anomaly-oriented manner, which is a different practice from mainstream IAD datasets. Anomaly categorization benefits ZSAS/IAD in terms of two aspects: (1) It enables us to appoint a certain anomaly class for ZSAS, which can be vital in some domains where certain classes of anomalies occur by a higher frequency. (2) It facilitates us to learn characteristics of a certain anomaly class by grouping them together. Here we categorize collected anomalies into 8 frequently-encountered industrial anomaly classes: Pits, Crack, Hole, Stain, Scratch, Orifice, Impurity, Abrasion. Having categorized collected anomalies, we ensure each anomaly image is assigned with a mask to indicate its anomaly pattern. For images from online search, reality photographing and datasets without masks, we annotate them by an annotation tool named EISeg [hao2022eiseg], while those anomaly masks from existing datasets can be directly fetched. In this way, our domain-agnostic anomaly dataset can be built in an efficient and extendible manner. More details of our dataset are shown in supplementary material.
@@ 3.3 CAI-guided Diffusion Mechanism (CDM)
With our new dataset, CAI has already been able to synthesize massive and diverse pseudo anomalies in a training-free manner. Although our dataset provides abundant real anomaly patterns, the quantity of anomaly patterns is still not infinite, which motivates us to break the quantity limit and further enrich anomaly patterns. Since CAI can synthesize highly authentic pseudo anomalies for a certain anomaly class, a natural idea is to feed them into a generative model as training data, so as to generate unlimited pseudo anomalies of this anomaly class. To this end, we propose CAI-guided Diffusion Mechanism (CDM), which is inspired by Latent Diffusion Model (LDM) [rombach2022high] and the recent FSAS solution [hu2024anomalydiffusion]. Specifically, for anomaly class $\mathcal{C}_{j}$ in a specific domain, CAI synthesizes $N^{\prime}$ pseudo anomaly images of anomaly class $\mathcal{C}_{j}$: $\mathcal{I}_{j}^{(sa)}=\{I_{i}^{(sa)}\}^{N^{\prime}}_{i=1}$, which are used as training data to guide anomaly generation. LDM is used as the generative model. To extract common anomaly features of $\mathcal{C}_{j}$, CDM uses Textual Inversion (TI) [DBLP:conf/iclr/GalAAPBCC23], which can learn those features by an embedding $e$ with very few samples (e.g. 3-5). The learned $e$ then serves as the condition that guides LDM to generate more pseudo anomalies of $\mathcal{C}_{j}$. To facilitate learning anomaly appearance and location separately, CDM follows [hu2024anomalydiffusion] to decouple $e$ into an anomaly embedding $e_{a}$ and a spatial embedding $e_{s}$: $e=\{e_{a},e_{s}\}$. To incorporate prior knowledge of anomalies into training, $e_{a}$ and $e_{s}$ are initialized as follows: For $e_{a}$, it is initialized by the word embeddings of a textual description about $I_{i}^{(sa)}$’s domain and anomaly class, e.g. “A photo of a capsule+crack”. As to $e_{s}$, we first build a location mask $y_{i}^{(loc)}$ to indicate the core location of injected anomaly pattern in the pseudo anomaly image $I_{i}^{(sa)}$:
$y_{i}^{(loc)}(l)=\left\{\begin{aligned} 1,&\qquad l\in\Omega(I^{(n)},l_{c},A^{\prime})\\ 0,&\qquad otherwise.\end{aligned}\right.$ (5)
where $I^{(n)},l_{c}$ and $A^{\prime}$ denote the target image, injection location and injected anomaly pattern of $I_{i}^{(sa)}$, while $\Omega(I^{(n)},l_{c},A^{\prime})$ refers to the region that $A^{\prime}$ occupies when it is centered at $l_{c}$ of $I^{(n)}$. Note that we use $y_{i}^{(loc)}$ rather than $y^{(sa)}$ in Eq. (4), so as to focus on the core anomaly pattern and ignore less important pixel changes caused by PE. Then, $y_{i}^{(loc)}$ is encoded into $e_{s}$ by a an spatial encoder $E_{s}$ (detailed in supplementary material). With initialized $e=\{e_{a},e_{s}\}$, we train a LDM $G_{a}$ to learn $e$ and $E_{s}$ [DBLP:conf/iclr/GalAAPBCC23]:
@@
where $E_{a}$ is the image encoder of $G_{a}$ that encodes $I_{i}^{(sa)}$ into a latent code $z_{i}$, $\epsilon\sim\mathcal{N}(0,1)$ is a Gaussian noise and $\epsilon_{\theta}$ denotes the noise estimated by $G_{a}$ parameterized by $\theta$ at time step $t$. Pixel-wise multiplication with $y_{i}^{(loc)}$ aims to make the model focus on learning the anomaly patterns in the injected region [hu2024anomalydiffusion]. Meanwhile, another LDM $G_{m}$ is trained as an anomaly mask generator $G_{m}$ by $y_{i}^{(loc)}$ and TI, which generates diverse anomaly masks to appoint the locations of generated pseudo anomalies. As to anomaly generation, $G_{m}$ first generates a new mask $y^{(g)}$ and then encodes it into $e^{(g)}_{s}$ by the trained spatial encoder $E_{s}$. Then, we fetch $e_{a}$ from the learned $e^{*}$ to denote anomaly appearance, and a condition embedding $e^{\prime}$ is yielded by $e^{\prime}=\{e_{a},e^{(g)}_{s}\}$. Finally, a new pseudo anomaly image can be generated by feeding $e^{\prime}$ and a target normal image $I^{(n)}{}^{\prime}$ into $G_{a}$. Besides, we leverage Adaptive Attention Re-weighting [hu2024anomalydiffusion] and blended diffusion process [avrahami2022blended] to ensure a fulfilled and smooth generation. Compared with the FSAS solution [hu2024anomalydiffusion], CDM’s biggest feature is to guide anomaly generation by CAI’s high-quality pseudo anomalies, rather than domain-specific anomalies. In this way, our ZSAS paradigm can also exploit merits of model-based methods, and CDM can serve as a powerful supplement to our model-free CAI method. Later experiments in Sec. 4.2 demonstrate CDM’s ability to extend pseudo anomalies under CAI’s guidance.

### v2

@@ 1 Introduction
To ease above problems, recent years have witnessed a surging interest in few-shot/zero-shot anomaly synthesis (FSAS/ZSAS), which aims to synthesize pseudo anomaly images for a specific industrial domain with few or zero domain-specific anomaly. A fair number of FSAS/ZSAS solutions have been proposed (reviewed in Sec. 2), while pseudo anomalies gradually occupy an indispensable role in many recent IAD solutions [zavrtanik2021draem, li2021cutpaste, schluter2022natural, zhang2024realnet, hu2024anomalydiffusion]. Since FSAS still requires accessing a few domain-specific anomalies in advance, FSAS suffers from the uncertainty or even unavailability of domain-specific anomalies. By contrast, ZSAS makes no assumption on the availability of domain-specific anomalies, thus enjoying better applicability to more diverse domains. Hence, we will focus on ZSAS in this paper. To maximize pseudo anomalies’ utility and facilitate them to be incorporated into IAD, a natural goal is to realize both authentic and simple ZSAS. Unfortunately, we observe that existing ZSAS solutions find it hard to balance authenticity and simplicity. Specifically, existing ZSAS solutions can be categorized into model-free solutions and model-based solutions: Model-free solutions (e.g. CutPaste [li2021cutpaste], DRAEM [zavrtanik2021draem]) play a dominant role in ZSAS. They sample patterns from given normal images or external datasets that are irrelevant to IAD, and directly inject those patterns into the target normal image to synthesize anomalies. In this way, such solutions enjoy a simple use without training an extra model. However, pseudo anomalies synthesized by model-free solutions typically lack authenticity, because their pseudo anomaly patterns are irrelevant to real anomalies. Besides, their authenticity is further diminished by injection artifacts (e.g. pixel discontinuity). By contrast, recent model-based solutions like RealNet [zhang2024realnet] place more emphasis on authenticity. They promote authenticity by training a deep generative model like Generative Adversarial Networks (GANs) [goodfellow2020generative] or Diffusion Models (DMs) [song2019generative], while the core idea is to re-direct the model’s generative process by some external factors (e.g. carefully-crafted perturbations) to generate pseudo anomalies. Although authenticity is improved, such solutions involve a cumbersome training process that requires heavy calculations and extra hardware. Meanwhile, re-directing the generative process is not straightforward, which often requires delicate model design and tuning. Therefore, model-based solutions lack simplicity due to their training-based nature, and their high cost limits their application to practical IAD.
@@
  • We build the very first domain-agnostic anomaly dataset (DAAD) within our best knowledge to supply CAI with rich real anomaly patterns from abundant cross-domain anomalies. It also benefits future ZSAS/IAD research.
  • We propose a CAI-guided Diffusion Mechanism, which can utilize CAI’s authentic pseudo anomaly images to perform unlimited anomaly synthesis. It enables us to further break the quantity limit of real anomalies in DAAD.
@@ 3.1 Cross-domain Anomaly Injection (CAI)
Given a set of cross-domain anomalies, we randomly sample a cross-domain anomaly image $I_{i}^{(ca)}\in\mathbb{R}^{H_{a}\times W_{a}\times 3}$ as source image, and its binary mask $y_{i}\in\{0,1\}^{H_{a}\times W_{a}}$ that indicates the area of its anomaly pattern $A_{i}=I_{i}^{(ca)}(y_{i})$, where $I_{i}^{(ca)}(y_{i})$ refers to the image pattern indexed by $y_{i}$’s non-zero elements. Then, we fetch a target image $I^{(n)}\in\mathbb{R}^{H_{n}\times W_{n}\times 3}$ from the normal training set of target domain and resize it to the target size for processing. Finally, CAI intends to inject $A_{i}$ into $I^{(n)}$ to synthesize a pseudo anomaly image $I^{(sa)}$. Concretely, CAI consists of four major stages: Source-target Matching, Multi-scale Anomaly Synthesis, Injection Location Selection and PE-based Anomaly Injection. The procedure of CAI is shown in Fig. 3 and Algo. 1, while we elaborate each stage below:
@@
Source-target Matching (STM). As the source image $I_{i}^{(ca)}$ and target image $I^{(n)}$ are from different domains, real anomaly pattern $A_{i}$ in $I_{i}^{(ca)}$ often requires resizing before injection to fit $I^{(n)}$. In this process, we need to avoid improper resizing, which makes injected anomaly patterns too trivial or too vague. To this end, we propose Source-target Matching. Specifically, we first extract a binary mask to indicate the foreground in $I^{(n)}$: $y^{(n)}=m(I^{(n)})$, where the function $m(\cdot)$ either returns the mask of the object by image matting [gatis_rembg_2022], or return an all-one mask for images without any object (e.g. texture). Afterwards, we calculate the scale of the anomaly pattern in $I_{i}^{(ca)}$ and the foreground in $I^{(n)}$, and define the original Source-Target Ratio (STR), $R$:
@@
Multi-scale Anomaly Synthesis (MAS). To promote the diversity in pseudo anomalies’ scale, CAI adopts a Multi-scale Anomaly Synthesis strategy. Given the target image, the number of large/medium/small pseudo anomalies, $N_{l}/N_{m}/N_{s}$, are explicitly set for anomaly synthesis. Meanwhile, MAS requires prioritizing synthesis of larger pseudo anomalies under the constraints of STM. For example, MAS prioritizes synthesizing large pseudo anomalies when given a large anomaly pattern, while medium or small pseudo anomalies are produced only when the required number of large pseudo anomaly is reached. Therefore, an anomaly pattern prefers synthesizing a pseudo anomaly of the same scale, but it can also synthesize pseudo anomalies in other scales. In this way, a rational $R^{\prime}$ is selected by STM and MAS (see line 6-23 of Algo. 1). In this paper, we synthesize $N_{l}=4,N_{m}=3,N_{s}=3$ pseudo anomalies with a target normal image, which is a relatively balanced ratio for performance and computational overhead.
@@
where $\partial\Omega(I^{(n)},l_{c},A^{\prime}_{i})$ denotes the border of region $\Omega(I^{(n)},l_{c},A^{\prime}_{i})$. Eq. (3) actually encourages $\Omega(I^{(n)},l_{c},A^{\prime}_{i})$ to share the pattern of the injected anomaly $A^{\prime}_{i}$ by minimizing their difference in gradients, while the constraint of Eq. (3) ensures a smooth transition at the region border. Eq. (3) is equivalent to solving a Poisson partial differential equation that subjects to Dirichlet border condition. Two modes can be selected in PE: For Normal mode, the injected pattern is solely fetched from the source image; For Mixed mode, the injected pattern is a mixture of source and target image pattern. We denote CAI using Normal and Mixed mode PE by CAI-N and CAI-M, respectively. Then, we filter out the injected anomaly pattern outside the foreground region by the foreground mask $y^{(n)}$. Considering all inputs, the whole PE-based injection process can be denoted by $I^{(sa)}=PEI(I^{(n)},y^{(n)},l_{c},A^{\prime}_{i})$.
@@
3.2 Domain-agnostic Anomaly Dataset (DAAD)
@@
One of CAI’s foundation is a real anomaly dataset that can provide abundant cross-domain anomalies. However, we notice that existing IAD datasets typically follow a domain-oriented fashion, i.e. they focus on collecting normal/abnormal images under certain domains, and anomalies are mainly categorized by their associated domain. For example, for the popular MvTec-AD dataset [bergmann2019mvtec], its primary categories are like “Bottle/Capsule/Tile…”. Unfortunately, this fashion suffers from two major drawbacks: (1) It hinders the collection of extensive real anomalies, as only anomalies in a certain domain are of interest. (2) It stresses the domain rather than anomalies. Nevertheless, CAI requires massive cross-domain anomalies, which makes existing IAD datasets short in quantity and diversity of real anomalies. This inspires us to build the first domain-agnostic anomaly dataset named DAAD, and we show its building procedure in Fig. 4 and detail its key steps below.
@@
Real Anomaly Collection. Real anomalies are actually abundant when the constraint of domain is removed. To collect domain-agnostic real anomalies, we exploit three sources that are all simple and affordable for the general public: (1) Online search: We feed IAD-related textual queries (e.g. “industry scratch”) into the search engine to obtain real anomaly images. (2) Reality photographing: We use frequently-seen portable photographing devices like digital cameras or cellphones to capture real anomalies. (3) Existing datasets: We exploit real anomalies from 9 existing IAD datasets (listed in supplementary material). By virtue of three sources above, we are able to collect about 6000 real anomaly images in a short period ($\sim$ 2 weeks) as the data foundation of our domain-agnostic anomaly dataset. Our anomaly collection is much more extensible when compared with classic IAD datasets, as DAAD welcomes any real anomaly from diverse domains. Note that if an existing dataset is used for evaluation, its data will be strictly excluded from ZSAS to prevent information leakage.
@@
Anomaly Categorization and Annotation. After collecting sufficient real anomalies, we categorize real cross-domain anomalies into several anomaly classes $\{\mathcal{C}_{1},\mathcal{C}_{2},...\}$ in an anomaly-oriented manner, which is a different practice from mainstream IAD datasets. Anomaly categorization benefits ZSAS/IAD in terms of two aspects: (1) It enables us to appoint a certain anomaly class for ZSAS, which can be vital in some domains where certain classes of anomalies occur by a higher frequency. (2) It facilitates us to learn characteristics of a certain anomaly class by grouping them together. Here we categorize collected anomalies into 8 frequently-encountered industrial anomaly classes: Pits, Crack, Hole, Stain, Scratch, Orifice, Impurity, Abrasion. Having categorized collected anomalies, we ensure each anomaly image is assigned with a mask to indicate its anomaly pattern. For images from online search, reality photographing and datasets without masks, we annotate them by an annotation tool named EISeg [hao2022eiseg], while those anomaly masks from existing datasets can be directly fetched. In this way, our domain-agnostic anomaly dataset can be built in an efficient and extendible manner. DAAD is open and its details are shown in supplementary material.
@@ 3.3 CAI-guided Diffusion Mechanism (CDM)
As shown above, our model-free CAI enables both efficient and effective ZSAS, but it relies on DAAD dataset. Although our DAAD dataset has already provided considerable real anomaly patterns, its further extension requires time and labor, which can be an impediment to CAI-based ZSAS. To break the quantity limit of real anomalies in DAAD, a natural alternative is to exploit high-quality pseudo anomalies yielded by CAI to train a generative model, so it can enable unlimited ZSAS and generate more diverse pseudo anomalies. To this end, we propose CAI-guided Diffusion Mechanism (CDM), which serve as a powerful model-based supplement alongside our model-free CAI solution. CDM is inspired by Latent Diffusion Model (LDM) [rombach2022high], Textual Inversion (TI) [DBLP:conf/iclr/GalAAPBCC23] and a recent FSAS solution [hu2024anomalydiffusion]. It is trained by the objective below:
@@
Due to page limit, see supplementary material for CDM’s details. Guided by CAI’s high-quality pseudo anomalies, CDM can generate more diverse pseudo anomalies with good authenticity (see Fig. 5), while experiments in Sec. 4.2 also show that pseudo anomalies generated by CDM can be readily applied to training a IAD model. Note that CAI is independent of CDM, as CDM is a supplement to CAI.